### v1

# Improving sampling of binding free energy differences between covalently bound ligands in alternate binding pockets using MT-REXEE

Abstract
The primary limitation for the application of alchemical free energy methods to a wider variety of complex molecular systems in achieving reasonable sampling. Flexible binding complexes often have high free energy barriers, which require prohibitively long simulations to sample sufficiently to obtain reliable free energy estimates. An example of such a system is the FabB–acyl chain complex, which features acyl chain substrates from 4 to 16 carbons in length. Previous experimental evidence supports that there are also at least two accessible binding pockets for the acyl chain bound to FabB. Sampling of both pockets would require prohibitively long MD simulations or the creation of complicated and nonobvious collective variables. With the multiple topology replica exchange of expanded ensemble (MT-REXEE) enhanced sampling approach, we can obtain highly efficient sampling between both pockets by adaptively growing and shrinking the chains in the simulation ensemble, allowing each simulation to visit chain lengths where transitions between the pockets do occur.
To enable this enhanced sampling for the chain growth problem, we demonstrate a new and improved conformational swapping selection algorithm for MT-REXEE, and validate this algorithm on a previously studied simultaneous calculation of multiple relative binding free energies. Using the new swapping approach gives significantly enhanced sampling even for this simpler problem, as demonstrated by faster convergence of free energy estimates.
This case study demonstrates the utility of MT-REXEE and its open-source implementation for systems that feature high free energy barriers for a subset of ligands of interest, demonstrating a valuable addition to the existing stable of enhanced sampling methods.
1 Introduction
Alchemical free energy (FE) calculations have become standard in many computational molecular researchers’ workflows, but their application is primarily limited to systems with comparatively little conformational flexibility 1, 2. Though the primary application of FE simulations has historically been small-molecule drug discovery 3, 4, their utility extends far beyond such binding free energy calculations. However, these less standard applications often involve systems that exhibit significant conformational flexibility, either in the protein, the ligand, or both, requiring prohibitively long simulations to converge FE estimates. In many cases, systems with significant flexibility can remain trapped in metastable states for longer than is reasonable to simulate them. These limitations have prompted the discovery of numerous enhanced sampling methods, including our own multiple topology replica exchange of expanded ensemble 5.
Protein and enzyme engineering is an important, emerging application for FE calculations in large and flexible systems 6, 7. Often, engineered protein systems involve multimeric complexes, which introduce flexible inter-protein interfaces. One such example is the enzymes within the fatty acid synthesis (FAS) cycle. This cycle is of interest for the industrial biosynthesis of a wide range of products, including bio-fuel precursors 8, 9. FAS involves enzymes that form either dimers or tetramers. The substrate for FAS enzymes is transported between each enzyme in the cycle by the acyl carrier protein (ACP) 10, 11, 12, 13, 14. Thus simulation of the ACP–enzyme complex involves multiple flexible binding interfaces.
The covalent intermediate for FAS is both biologically relevant and a more straightforward state to quantify sampling enhancements than the ACP–enzyme complex. The covalent intermediate forms after the acyl chain substrate has been passed to the CYS residue on the respective FAS enzyme. The covalent intermediate state for FAS enzymes suffers from poor sampling in MD simulations due to the presence of a large, flexible substrate whose conformational flexibility is restricted by the covalent bonding to the protein and a relatively narrow binding pocket.
These same sampling challenges also occur when applying FE simulations to covalent drug development 15, 4, 16, 17 as covalent drugs have a similar covalent intermediate. Covalent drug development is an emerging field given the increased potency and lifetime within the system. However, enhancing the sampling of AFE methods by capturing multiple binding and unbinding events as alchemical parameters change to minimize the configurational free energy barriers is not possible for covalent ligands using molecular mechanics simulations, as covalent bonds cannot be formed or broken. FE calculations with covalent ligands, both in the enzyme engineering and covalent drug cases, thus require alternate methods of achieving enhanced sampling to improve the accuracy and utility of computational methods.
Existing enhanced sampling methods are insufficient for the flexible covalent ligands in the FAS system. Enhanced sampling methods that purely use the alchemical variable present in free energy calculations, $\lambda$, to enhance sampling include Hamiltonian replica exchange 18 and expanded ensemble 19, 20. These methods to enhance sampling require that either the end states or the intermediate states have significantly different free energy barriers in configuration space. For the FAS examples proposed earlier, we are interested in the free energy of binding as a function of substrate chain length, which can vary by 2 carbon intervals from 4 to 16 carbons. Performing a single alchemical transformation from 4 to 16 carbons would not give information about the intermediate chain lengths, and performing iterative transformations would not result in sufficiently low FE barriers between alternate bound configurations at long chain lengths to converge sampling.
Collective variable (CV) dependent methods, including metadynamics 21, 22, umbrella sampling 23, 24, and weighted ensemble approaches 25, represent another approach to enhanced sampling. However, they require significant prior knowledge about the conformational changes that enable transition between trapped metastable states. For the FAS examples, although we know that longer chains must be able to move between pockets, the number of degrees of freedom involved makes it challenging to identify relevant CVs that can describe such a transition involving all of the atoms. Intermediate motions along the proposed CV likely sample high-energy configurations, which decrease sampling efficiency. Even with an approximate path, these CV-dependent methods require prior knowledge of where any alternate pockets may be or significant trials to derive a more complete CV.
We previously presented the enhanced sampling method, multiple topology replica exchange of expanded ensemble (MT-REXEE), which specifically targets enhancing sampling in flexible receptor-ligand binding systems 5. In this work, we present improvements made to the MT-REXEE method through the introduction of a novel swapping algorithm, random range swapping, which increases the efficiency of the calculations and improves FE estimate convergence. This method works by enabling trajectory frames other than the last one in the iteration to be used as potential swappable states rather than utilizing redundant end-states. We also apply MT-REXEE to the difficult sampling problem, calculating the acyl chain length dependence on the binding FE for the FabB-acyl complex. This complex demonstrates the ability of MT-REXEE to significantly enhance conformational sampling in bound systems by coupling free energy simulations together configurationally.
The FabB-acyl complex represents an excellent test case demonstrating how MT-REXEE can accelerate sampling without prior knowledge of the behavior of the system. Previous experimental evidence revealed that the I108M mutation in FabF significantly reduces synthesis of acyl chain products longer than 8 carbons in length 26. The analogous mutation G107M in FabB, however, did not produce the same reduction of activity on medium chain substrates 26. We hypothesize that an alternate binding pocket is accessible in FabB, which enables FabB to continue synthesis even in the event that the canonical binding pocket is blocked by the introduced MET residue. Here we demonstrate the ability of MT-REXEE to sample both the canonical binding pocket (pocket A) and the alternate binding pocket (pocket B) in both WT and G107M FabB variants, without biasing the simulation towards the alternate binding pocket or even identifying it beforehand. With MT-REXEE, not only do we access both pockets within the same simulation, but we can quantify the FE penalty of the G107M mutation in pocket A and the relative stability between of pockets A and B.
Sampling of both distinct binding pockets is not possible using standard MD or independent expanded ensemble simulations within reasonable simulation lengths (<<500 ns) for substrates longer than 4 carbons. But we can leverage the increased conformational flexibility of the short-chain substrates through these conformational swaps to increase sampling in long-chain simulations. We present this method as a suitable method for overcoming high configurational free energy barriers which are present for only a subset of ligands being evaluated. We also demonstrate how MT-REXEE simulations can produce both configurational and alchemical free energy values for all evaluated end points and between two or more distinct binding pockets.
2 Theory: Separating Alchemical and Conformational Free Energies
One application of the MT-REXEE method is to reconstruct the binding FE dependence on both ligand identity and bound conformation for systems that feature two or more distinct binding locations. We apply the MT-REXEE method to the FabB acyl covalent intermediate complex, as this system features a wide variety of ligand substrates (acyl chain substrates vary from 4 to 16 carbons) and is expected to sample two distinct acyl chain binding locations (A and B). For this system, each MT-REXEE simulation performs one alchemical transformation and can feature two or more configurational states (Figure 1A). MT-REXEE enables conformational exchange between each transformation, enabling more comprehensive sampling of the two or more conformational states. From the set of MT-REXEE simulations, we obtain the dependence of binding FE on both the substrate and the binding conformation (Figure 1B).
Figure 1: (A) A set of MT-REXEE simulations which perform transformations of the acyl substrate from C2 to C4, C4 to C6, and C6 to C8. Each simulation, visiting a range of alchemical states $\lambda$ in [0,1], has configurations that should (in the limit of sufficient sampling) visit both pocket A (orange) and pocket B (brown). The dashed boxes indicate states in each individual simulated transformation, and the dark blue arrows show where configurational swaps can occur, at the alchemical end states. (B) When we separate the transformations by end state and configurational state, we can estimate each of the FE differences we compute from a MT-REXEE simulation. We compute the alchemical FE difference for each transformation independently in each pocket (light blue) as well as the configurational FE difference at at least one end state (pink). This enables us to compute the relative FE difference between any combination of alchemical end state and configurational state.
First, we demonstrate that we can partition the conformational space to calculate the alchemical free energy difference between two end-points independently for each configurational state. This calculation will provide the values for $\Delta\Delta G_{n,A}$ and $\Delta\Delta G_{n,B}$ (light blue arrows) in figure 1B, where $n$ indicates the transformation. We start with the alchemical Gibbs free energy difference given by equation 1 where $U_{i}$ and $U_{j}$ are the potential energies as a function of $\vec{q}$ the coordinates and momenta for two alchemical states and $\Gamma_{i}$ and $\Gamma_{j}$ denote the phase space volumes of $\vec{q}$ over which we sample, though for most alchemical free energy simulations $\Gamma_{i}=\Gamma_{j}$.
$\Delta G_{ij}=-k_{B}T\ln{\frac{Q_{j}}{Q_{i}}}=-k_{B}T\ln{\frac{\int_{\Gamma_{j}}e^{-\frac{U_{j}(\vec{q})}{k_{B}T}}d\vec{q}}{\int_{\Gamma_{i}}e^{-\frac{U_{i}(\vec{q})}{k_{B}T}}d\vec{q}}}$ (1)
The total phase space can be decomposed into the space consistent with sets of conformations in pocket A and B, such that $\Gamma_{i}=\Gamma_{i,A}+\Gamma_{i,B}$. Here we are interested not in the overall relative free energy difference between the two alchemical states ($i$ and $j$) but rather the relative free energy difference between the two alchemical states within the restricted phase space consistent with configurational state A and thus equation 1 becomes 2. $\Delta G_{ij,A}$ is the free energy difference of performing the alchemical transformation from state i to j while in pocket A, and $\Delta G_{ij,B}$ would give the same transformation but in pocket B.
$\Delta G_{ij,A}=-k_{B}T\ln{\frac{Q_{j,A}}{Q_{i,A}}}=-k_{B}T\ln{\frac{\int_{\Gamma_{j,A}}e^{-\frac{U_{j}(\vec{q})}{k_{B}T}}d\vec{q}}{\int_{\Gamma_{i,A}}e^{-\frac{U_{i}(\vec{q})}{k_{B}T}}d\vec{q}}}$ (2)
The alchemical free energy difference within conformational space $A$ and $B$ can be calculated by dividing the trajectory frame by frame into two non-continuous trajectories corresponding to spaces $A$ and $B$ based on the configurational state of each frame. We can then perform two separate MBAR calculations (Figure 6) with the samples from each trajectory. The two partition functions are additive, but the two alchemical free energies extracted from each transformation are not additive, which we expand upon in section Non-additive nature of $\Delta A_{ij,A}$ and $\Delta A_{ij,B}$.
There may be concern that the dimensionality of the conformational space defining A and B will vary between transformations as each transformation is performed with different numbers of dummy particles. At this point, it is important to note that this method is being used purely for differences in free energies. This means that an equivalent transformation, between states $i$ and $j$, is performed in solvent as well as in complex with the protein. Since the calculation of $\Delta\Delta G$ is independent for each transformation and we maintain only three bonded interactions between the dummy and ”real” atoms during the transformation, Boresch shows that the dummy atom contributions cancel, assuming equal dummy atom treatment in both solvent and protein complex 27. We could rectify the issue of dimensionality differences in partition space through the addition of dummy atoms to all transformations, such that all atoms necessary for all transformations are present, but many remain in the dummy state for both end-points of the FE transformation. Boresch shows that the configurational free energy (Z) is given by equation 3.
$Z$ $=Z_{J}Z_{vib}Z_{PMF}$ (3)
$Z$ $=Z_{J,D}Z_{J}Z_{vib,D}Z_{vib}Z_{PMF,D}Z_{PMF}$ (4)
In this equation, $Z_{J}$ is the contribution from the Jacobian, $Z_{vib}$ is the contribution from the vibrational motions, and $Z_{PMF}$ is the contribution from the potential of mean force. In the same way that we separate the contribution for the dummy atoms that change alchemical state during the transformation (Equation 4), we can use the same logic to add a second class of dummy atoms that remain in the decoupled alchemical state. Their contribution would be statistically identical for both sides of the free energy cycle and would thus cancel in the same manner.
In this application, the distinction between pocket A and B exists solely in the protein complex conformation and thus in solvent $\Delta G_{ij,A,solv}=\Delta G_{ij,B,solv}$ but this is not necessarily the case so the two will be left as independent variables. Thus the relative binding free energy between alchemical states $i$ and $j$ ($\Delta\Delta G_{ij,A}$) can be given by equation 5 (see Figure 7 for the full thermodynamic cycle). For simplicity, when dealing with multiple subsequent transformations, we instead use the notation $\Delta\Delta G_{n,A}$ where n denotes the transformations in order from the base transformation.
$\Delta\Delta G_{n,A}=\Delta G_{n,A,complx}-\Delta G_{n,A,solv}$ (5)
Finally, we need to compute a configurational FE difference between pockets A and B, i.e., $\Delta G_{conf,0}$ (pink arrow) in Figure 1B. In standard MD simulations, which sample from the “native” Boltzmann distribution with no conformational biases, the relative probability of sampling within the conformational space of pocket A versus pocket B is given by the integrated Boltzmann factor over the phase spaces volumes corresponding to the configurations; the configurations where to the chain occupies pocket A versus the set of configurations where the chain occupies in pocket B. for Boltzmann sampling, then this becomes simply the ratio of time (i.e. number of frames) spent in each configurational state. This is shown in Equation 6 where $p_{A}$ and $p_{B}$ and the probability of sampling within the conformational space of pocket A and pocket B respectively, $\beta=k_{B}T$, $k_{B}$ is the Boltzmann factor, $T$ is the temperature, and $\Gamma_{A}$ and $\Gamma_{B}$ are the sets of configurations corresponding to the chain being in pocket A and B respectively.
$\frac{p_{A}}{p_{B}}$ $=$ $\frac{Q_{A}}{Q_{B}}=\frac{\int_{\Gamma_{A}}e^{-\beta U(\vec{x})}d\vec{x}}{\int_{\Gamma_{B}}e^{-\beta U(\vec{x})}d\vec{x}}$ (6)
$\Delta G_{conf,AB}$ $=$ $k_{B}T\ln\frac{p_{A}}{p_{B}}$ (7)
In this application, we use our EE alchemical free-energy simulations to compute the configurational free energy rather than standard MD, which means we cannot extract the configurational free energy from the entire simulation trajectory to attempt to compute probabilities with the simulation or pocket A and pocket B, as it features non-physical intermediate states. Instead, we can extract a subset of the sampled configurations from the physically meaningful end states ($\lambda=0$), and we can apply equation 6 to determine the configurational free energy difference. We require only one configurational FE difference between states. This choice should be the state for which the kinetics of exchange between configurational states is the fastest. In this application, this is the C2 acyl chain end state. This state had statistically equivalent pocket sampling to that of the independent MD simulations (see Figure 3B), and the uncertainty in this step is less than the uncertainty in the alchemical process 5. Thus we can use the relative probability of the C2 chain sampling either pocket A or B with equation 7 to compute the configuration FE difference $\Delta G_{conf,0}$
This configurational FE difference ($\Delta G_{0,conf}$) along with the alchemical free energy differences in each transformation ($\Delta\Delta G_{n,A}$ and $\Delta\Delta G_{n,B}$) can then be used to compute the FE difference between any end state in any configurational state. Thus, the configurational free energy difference for an end point after $x$ transformations is given by 8 (Figure 15). This relationship will change if the choice of flexible conformational state used to compute the configurational FE difference is different, but we can use any sum of paths which connect the two relevant end states.
$\Delta G_{x,conf}=\Delta G_{0,conf}+\sum_{n=0}^{x}\Delta\Delta G_{n,A}-\Delta\Delta G_{n,B}$ (8)
3 Methods
3.1 System Set-up
We used two systems to test this approach, MUP1-thiazole and FabB-acyl complexes described in the introduction. The MUP1-thiazole complex features a cyclic ligand bound to a relatively rigid $\beta$-barrel protein domain for which convergence of free estimates requires easily accessible computational resources 28. We used the FabB-acyl complex allows us to demonstrate the conformational sampling enhancement of the MT-REXEE approach as it features high free energy barriers for acyl chain rearrangement, which as described above are difficult to sample with existing methods. The protein-thiazole complexes were derived from the crystal structure (PDB 1IO6). The FabB covalent intermediate complexes were derived from the crystal structure of WT FabB with a covalently bound C12 acyl chain (PDB 1EK4). MODELLER was then used to apply the G107M mutation. All small molecules were parameterized using GAFF2 29 and the protein was parameterized using the Amber ff14SB force field30. The relative binding free energy topologies for all systems were created using PMX 31 followed by post-processing to remove all dihedral energies between dummy and real atoms within the hybrid topologies. The treatment of dummy atoms in this study follows the practices presented in earlier works to minimize any effects on the partition function that are not canceled out by alchemical transformations in different environments  32, 33. The unit cell boxes for all simulations conducted in solvent were constructed to maintain a minimum distance of 1 nm between the molecule and the periodic boundary and then filled with TIP3P 34 solvent. All simulations were run in a modified version of GROMACS 2022.5 35, which can be found as a branch on the official GROMACS GitLab repository; all code changes in this branch have been merged into the main repository as of GROMACS 2025.0. This modification allows the acceptance of the MDP parameter init_histogram_counts, which is crucial for performing Wang-Landau weight equilibration across iterative trajectories.
[…]
We initiated the FabB covalent intermediate complex with the acyl chain in two different, non-overlapping pockets, which are referred to as pocket A and pocket B (Figure 3A). Equilibration was performed for the FabB–acyl complex in pocket A with 50 ns using standard MD at $\lambda=0$. From this simulation, three random frames were selected in the first, second, and last third of the trajectory as the initial configuration for each of the three replicas. We then run MT-REXEE simulations for the FabB–acyl complex in pocket A. These simulations sample both pocket A and pocket B. We extract one frame in which each complex samples pocket B from these MT-REXEE simulations. We then ran 50 ns of standard MD on each of these pocket B configurations at $\lambda=0$ and again extracted a random frame from the first, second, and last third of the MD trajectories as the initial configurations.
3.2 Free Energy Protocol
We performed both EE and MT-REXEE simulations in triplicate for each system. The EE settings used were identical for both the MT-REXEE and independent EE simulations with a Wang-Landau scale of 0.82 and a Wang-Landau ratio of 0.85 for the equilibration for the FabB system, and 0.8 was used for both settings for the MUP1 inhibitor binding free energy. These parameters were tuned to ensure relatively even sampling between $\lambda$ intermediate states following weight equilibration. Slower weight equilibration through the increasing of Wang-Landau scale and ratio parameters extends equilibration times but produces more even $\lambda$ state sampling once equilibrated weights are obtained. The weight equilibration phase was ended when the Wang-Landau $\delta$ was less than 0.001 for both systems, a cutoff determined from previous experience. The MUP1 inhibitors were run for 30 ns and 50 ns after equilibration for the solvent and protein complex simulations, respectively. We simulated the FabB complexes in the covalent intermediate state, but since the ligand was previously covalently bound to the protein ACP, the“solvent” simulation state was defined as ACP with the acyl chain in solvent. The FabB systems were run for 50 ns and 100 ns for the solvent and enzyme-bound states, respectively. All free energy estimates were calculated using the MBAR implementation in alchemlyb 38.
The primary feature added to the MT-REXEE package in this study is the use of a new swapping criterion between simulations. In the previously developed exhaustive approach, swaps are only attempted for the last frame of each iteration. This new swapping criterion is named random range (RR), and rather than only allowing swaps between trajectories in the last frame of an iteration, we instead allow swaps at random compatible frames (i.e. endpoints of two neighboring simulations) in a defined portion of the trajectory. We used the last 50% of the trajectory to avoid swaps being highly correlated to the previous iteration. Regardless of the location of the trajectory used for swapping, all data from each iteration is used to compute free energy estimates and any additional analysis; only trajectory swapping and continuation use the randomly sampled exchange-compatible frame. For the MUP1 system, we used redundant end states (5 intermediate $\lambda$ states as well as 5 states each at $\lambda$=0 and $\lambda$=1) in combination with the exhaustive swapping method as we did in our previous MT-REXEE paper 5 to compare to the random range swapping approach without redundant end states. For the FabB–acyl systems, we only ran MT-REXEE simulations with the RR method and with single $\lambda$ end-states.
3.3 Configurational Analysis
X-ray crystallography structures of the FabB–acyl-ACP complex have previously identified two distinct binding tunnels for the acyl chain. We predict that these distinct binding tunnels are also occupied in the covalent intermediate states, but there are no experimental structures defining this alternate binding pocket, which we term pocket B, in the covalent intermediate state. We used the MT-REXEE simulations with the acyl chain initialized in the canonical binding pocket, which we term pocket A. We then visually labeled a subset of trajectory frames (12–15) for all chain lengths as being in either pocket A or B in order to then determine a quantitative measurement. We determined that the distance between carbon 4 in the acyl chain and residue THR300 could be used to differentiate the binding pockets (Figure 8A). We evaluated several distance thresholds between pairs of atoms in the acyl chain and FabB as well as acyl internal dihedrals as metrics for pocket determination, but the distance between C4 and THR300 metric was determined to have the least overlap in values between pocket A and pocket B. We plotted the distance distribution of these residues in EE simulations which were visually verified to have all frames within either pocket A or B, as well as for frames from our MT-REXEE simulations which sample both pockets to determine the threshold of 0.85 nm (Figure 8B). Several conformations for longer chains (8 or more carbons) were identified as belonging in pocket A or B by this criterion, but did not match existing known structures for the pockets. Thus, we also applied a chain compaction criterion that the distance between carbon 2 and 8 should be shorter than 0.6 nm in pocket B and longer than 0.75 nm in pocket A determined from available crystal structures. Any trajectory frames in which this distance to THR300 is greater than 0.85 nm and the compaction distance is greater than 0.75 nm are considered in pocket A. Any frames with a distance to THR300 less than 0.85 nm and a compaction distance of less than 0.6 nm is considered in pocket B. Any frames for which neither pocket criteria were satisfied were discarded from free energy analysis, as they are in neither pocket (<<2% of trajectories).
4 Results
4.1 Random range swapping improves round trip time and statistical convergence compared to previous exchange methods
In this section, we compare the results for running binding free energy calculations for a set of small molecule ligands with the protein MUP1 using both the previous introduced exhaustive method and the random range (RR) method (Figure 2A–B). We ran this comparison to determine any changes in performance and ensure that this new algorithm for selecting swappable states does not introduce additional error. The RR method no longer requires redundant end states. This eliminates the increase in Wang-Landau weight equilibration times observed in our earlier work 5. The result of this improvement is that MT-REXEE with the RR exchange method (MT-REXEE RR) has equivalent or lower total Wang-Landau weight equilibration times compared to independent EE simulations (Figure 9). We note that weight equilibration times are calculated using the full iteration in which the weights reach the equilibration criteria, even if conformational exchanges occur at a frame earlier than the final frame of the iteration. Only the E-F transformation had a potentially statistically significant difference between the weight equilibration times using EE and MT-REXEE RR ($p=0.083$). We hypothesize that the decrease in weight equilibration times for the E-F transformation is likely due to the improved conformational space sampling using RR. This transformation is the most chemically distinct of all the MUP1 ligands evaluated in this study and for this reason, benefits most from the increased conformational sampling.
Figure 2: Random range swapping yields statistically equivalent free energies while significantly decreasing the round trip time through alchemical states. (A) A diagram of the sequence of ligand transformations used in binding free energy estimates with the protein MUP1 to validate the random range swapping scheme. The curved arrows show where MT-REXEE swaps occur. (B) The structure of MUP1 in complex with ligand A. (C) We compare the binding free energy estimates using $E$, MT-REXEE with the exhaustive swapping scheme, and MT-REXEE with the RR swapping scheme. We observe no significant differences between the free energy estimates for all three methods as the values are all within statistical error of one another. (D) We observe a significant (p<<0.001) decrease in the round-trip times from 10-20 ns to 3-5 ns for both solvent and MUP1 complex simulations using the RR swapping scheme compared to the previous exhaustive scheme.
The binding free energy estimates are statistically equivalent for EE and MT-REXEE with both exhaustive and RR swapping methods, but various convergence metrics indicate that less simulation time is necessary for equal accuracy with RR (Figure 2C). We measured the convergence of these free energy estimates both along the length of individual trajectories as well as convergence between trajectories initiated with significantly different initial configurations. Importantly, we observe that for both solvent and, even more significantly for the protein complex simulations, the free energy estimates converge significantly more quickly, resulting in closer estimates between the estimates for the first and last half of the trajectories for MT-REXEE RR compared to both EE and the MT-REXEE simulations using the exhaustive method (Figure 10A–B). We also notice a significant decrease in the standard error of the estimate between replicates of between 40 and 160% compared to EE (Figure 10C). These independent replicates were initiated with the same initial configurations extracted from standard MD trajectories separated by a minimum of 20 ns. The exchanges between distinct alchemical transformations in the MT-REXEE method enable increased sampling of the conformational space, which produces faster convergence. The standard error is also reduced by between 16-90% when using the exhaustive swapping method with redundant end states compared to EE. This is likely because the increase in conformational sampling is balanced by the decreased sampling of intermediate $\lambda$ states due to the addition of redundant end states. This results in decreased conformational sampling than MT-REXEE RR, but still faster conformational sampling than EE. This is consistent with the observed significant decrease in the round-trip times for the MT-REXEE RR method compared to the exhaustive swapping method (Figure 2D). We expect the over 4-fold decrease in round-trip time is primarily responsible for the faster increased conformational sampling in MT-REXEE RR compared to the MT-REXEE exhaustive method.
4.2 Enhanced Sampling of FabB with MT-REXEE
We apply the MT-REXEE method to an example system in which there is a known sampling limitation. Prior evidence suggests that in the FabB-acyl-ACP complex the acyl chain samples two distinct binding tunnels, pocket A and pocket B. Here we examine the FabB-acyl intermediate complex, after the transfer of the acyl chain from ACP to FabB through the formation of a covalent bond with CYS163 on FabB (Figure 3A). Testing sampling with the FabB-acyl complex simplifies assessing the sampling compared to the FabB-acyl-ACP complex, as the motions of ACP produce additional sampling issues as the FabB–ACP interface is large and flexible. Using the FabB-acyl complex allows us to focus on only the motions of the acyl chain. The FabB-acyl complex is also the well-known intermediate conformation before the condensation reaction with malonyl-ACP and the covalent intermediate to extend the acyl chain by two carbons, making it a biologically important state 10.
We did not know whether the FabB-acyl covalent intermediate step would sample pocket B as was observed with the FabB–acyl-ACP complex, but previous mutational data did support sufficiently stable binding to enable production of acyl chains up to 16 carbons in length via FabB with a mutation which blocked pocket A through steric hindrance. We verified that the transition between the pocket A and B is too slow to sample with direct simulation using standard MD alone within reasonable simulation lengths (500 ns; Figure 3B). Therefore, this study uses MT-REXEE with random range swapping to evaluate the relative stability of 2, 4, 6, 8, 10, 12, 14, and 16 carbon length acyl chain substrates covalently bound to FabB. We only attach one substrate in a given simulation even though FabB is simulated in its native dimer conformation. This is motivated by existing evidence which suggests there may be negative cooperation between the two dimer members and it is thus more likely that only one reacts at a time.
We performed the 500 ns standard MD simulations for FabB-acyl complexes with acyl substrates from 2–16 carbons in length, seven simulations in all. All simulations were initiated with the same initial configuration as the corresponding MT-REXEE simulation in pocket A, with the removal of dummy atoms required for MT-REXEE. We ran three replicates with distinct initial configurations for all chain lengths (see Methods). The corresponding MT-REXEE simulations were run for only 100 ns (in parallel) for each transformation of chain length. We then analyzed the trajectories to determine the occupancy of pocket A, B, or neither for each chain length substrate. For MT-REXEE simulations, we specifically compared the pocket occupancy for the end states and not for intermediate $\lambda$ states, as these are irrelevant to the physical occupancies.
Figure 3: (A) The binding pocket of the FabB–acyl covalent intermediate is shown with a 12 carbon acyl substrate in both pocket A (orange) and pocket B (brown) in the context of the overall protein context. The structures shown are the centroids from MT-REXEE simulations of the C10-12 transformation for $\lambda=1$. (B) Occupancy of pocket B as a function of substrate chain length using both standard MD simulations and MT-REXEE enhanced sampling as described in section $3.$ Standard MD is not able to access pocket B within 500 ns for any substrate longer than 4 carbons, but we readily see sampling in pocket B for all chain lengths within only 100 ns of MT-REXEE simulation. (C) A close-up image of the binding pocket in (A) highlighting several residues on both the catalytically active monomer and adjacent monomer that restrict the acyl chain’s conformational flexibility and prevent interconversion during MD simulations of at least 500 ns.
MT-REXEE and standard MD show statistically equivalent sampling of pocket B only for the 2 carbon acyl chain substrate (Figure 3B). We found there is minimal sampling of pocket B for the 4 carbon substrate in standard MD and for all longer substrates there is no sampling of pocket B in standard MD started from pocket A. MT-REXEE thus enables significant sampling in pocket B for all chain lengths studied while utilizing only 20% of the trajectory length as the standard MD simulations.
We hypothesize that the inability of standard MD to sample both pockets is due to the presence of several acyl-chain-stabilizing interactions present for all acyl chain lengths 4 carbons or longer (Figure 3C). For both MD and MT-REXEE simulations, the 2-carbon acyl chain substrates are readily able to sample both pockets A and B with relatively high frequency. The presence of the conformational swaps between the 2 carbon state and longer chain substrates allows for conformational sampling of both pockets at all chain lengths. MT-REXEE enables simulation trajectories to sample pocket B when simulations are initiated in the canonical binding pocket A, thus requiring no prior knowledge of the presence or location of the alternate pocket. In contrast, existing knowledge of both the location of pocket B and a reasonable transition path would be necessary for employing a configurational variable dependent method. The lack of directly applied configurational bias along a known CV, like that used in metadynamics, also provides increased confidence that these alternate conformations are not artifacts of the enhanced sampling method itself.
In MT-REXEE simulations, switching between pockets is primarily limited by the round-trip time between alchemical simulations of different chain lengths, as we can only generally only switch between pockets in simulations of the shortest chain. Additionally, we want to emphasize that the success of the MT-REXEE method is heavily correlated to the suitability of the applied equilibrated EE weights. An increase in round-trip time correlates with an increased imbalance between $\lambda$ state sampling within the simulations (Figure 11). The imbalance in $\lambda$ state sampling can be minimized by adjusting Wang-Landau equilibration parameters. Optimization of these parameters is system-specific. Further optimization of these parameters would likely further minimize this effect, but would also increase weight equilibration times. This necessitates a balance between the efficiency of weight equilibration and the accuracy of equilibrated weights.
4.3 Configurational and Alchemical Free Energy
This sampling of both pockets during the same simulation allows us to achieve two goals: Estimating the relative free energy as a function of chain length independently in each pocket and estimating the configurational free energy difference for a given chain length substrate in pocket A versus pocket B. The solvent reference state for all alchemical binding free energies is the acyl chain attached to ACP in solvent as described in the Methods section (Figure 12).
We first ensured that the overlap between adjacent $\lambda$ states was sufficient for all transformations both in solvent (Figure 13) and in complex (Figure 14). We used the same $\lambda$ state spacing for both the EE and MT-REXEE simulations, as well as the same initial configurations for all replicas.
The EE simulations were initiated with the same chain conformations as the MT-REXEE simulations in pocket A. Since the EE transformations initialized in pocket A are only able to sample pocket A, we can only make comparisons to the alchemical free energy transformations in pocket A. We observe a significant difference between the free energy estimate for the C4 to C6, C6 to C8, and C12 to C14 transformations when using MT-REXEE compared to EE (Figure 4A). It is not unexpected that some transformations would produce significantly different FE estimates as even within a given pocket the configurational space sampled with each of the two methods is expected to vary greatly. This is less likely to affect the known flexible transformations like C2 to C4, which would be expected to readily sample the conformational space in both methods. The degree of sampling with a given binding pocket is also less likely to affect the longest substrates which have fewer accessible conformations which fit the full acyl chain. Overall, these results are consistent with the significantly different conformational sampling expected with MT-REXEE compared to EE. We would worry about MT-REXEE introducing additional error into the calculations producing these differences; however, the fact that the C2 to C4 transformation, the one transformation that we would expect EE alone to sample effectively, is statistically identical between the two methods suggests no errors are introduced.
Figure 4: The convergence of binding free energy simulations initiated in pockets A and B demonstrates MT-REXEE can increase sampling of the conformational space. (A) Comparison of estimates from individual independent EE calculations with estimates of the same transformations in MT-REXEE simulations. MT-REXEE’s significantly improved sampling leads to moderate differences between the two methods, with MT-REXEE expected to be more accurate. Estimates for binding free energy using MT-REXEE with the initial acyl chain configurations in either pocket A (brown) or pocket B (orange) for the calculation of the RFE in pocket A (B) and pocket B (C), showing independence of starting configuration. (D) Cumulative relative binding free energy can be obtained for all chain lengths relative to a C2 chain for both pockets using simulations initiated in both pockets A and B (Figure 15). Thus, from a single set of MT-REXEE simulations, we can obtain the full chain length and pocket dependence on binding FE. For all figures, we plot the mean of n=3 (n=6 for D) replicas and error bars represent the standard error of the mean.
We also performed MT-REXEE simulations initiated in pocket B to assess the convergence of the free energy estimates between the pockets, when the simulations were started in pocket A versus started in pocket B. After 100 ns of MT-REXEE simulation, the estimates were within statistical error regardless of whether they started in pocket A (Figure 4B) or pocket B (Figure 4C) to within statistical error, with the possible exception of the C14 to C16 transformation initialized in pocket A. A single simulation of this transformation which was initiated in pocket B experienced significantly less sampling in pocket A than the other two replicas. This contributes significantly to this deviation. As the simulation get longer, additional metastable states may emerge, so increase in sampling difficulty for very long chains is not unexpected.
Having established convergence between the two sets of initial configurations, all future presented data will use both sets of simulations regardless of initialization in pocket A or B. From these combined simulations, we compute the cumulative free energy curve with all estimates relative to the binding free energy of a C2 acyl chain in pocket A (Figure 4D).
Our results indicate that for acyl chains of 10 carbons or fewer, binding to pocket A in WT FabB is significantly more stable ($p<0.05$). For chains longer than 10 carbons pocket B is either statistically equivalent or significantly more stable than pocket A. We cannot perform a quantitative comparison to experimental data as there is no available experimental data on binding free energy for the covalent intermediate or on the relative occupancy of pocket A or pocket B. Importantly, these calculations also do not consider the relative stability of the ACP-complex configuration which would be necessary for direct quantitative comparison to experimental biding free energies. Given the complexity of interpreting the accuracy of our free energy estimates from the available experimental data for WT FabB alone, we compare the free energy surface between the WT and G107M variant of FabB.
The G107M variant is known to prevent the synthesis of fatty acids past 8 carbons in length in FabF, an enzyme isoform which does not have pocket B. The available experimental data on activity, along with the location of the mutations within the binding pocket, suggest that binding to pocket A is likely to be destabilized for substrates longer than 8 carbons in length (Figure 5A). We applied the MT-REXEE approach to measure binding free energy estimates for all chain lengths in both pocket A and pocket B to determine if this mutation affects the stability of the covalent intermediate state for chain lengths greater than 8 and to quantify that effect.
We compared the relative binding free energy values for all chain lengths in pocket A and pocket B between the WT and G107M FabB variants (Figure 5B–C). We observed that the G107M mutation appears to have a stabilizing affect on the complex for C2–6 chain lengths, as shown by the negative free energy difference between the mutations, shown in Figure 5D. The larger residue at position 107 is able to form additional non-bonded interactions within the pocket as well as with the short acyl chains which may explain this stabilizing affect. However, at the C6 to C8 transformation we see a significant destabilization of pocket A in the G107M variant which is not present for the WT as shown by the positive free energy difference in figure 5D. This difference is particularly evident for pocket A but not pocket B as shown by the difference in relative binding free energy between pocket A and pocket B ($\Delta G_{A}-\Delta G_{B}$). In the WT variant, chain growth into pocket A is more stable than pocket B until 6 carbons (subfigure D), the two pockets are statistically identical until 10 carbons, and then for 12 carbons or greater pocket B is more stable for each chain growth step into the pocket. However, in the G107M variant, pocket B is significantly more stable while growing the chain from 6 to 10 carbons (Figure 5D). This stark difference in relative binding free energies between pockets A and B coincides with the chain length at which binding in pocket A is predicted to be destabilized based on experimental evidence.
The focus of this work is on establishing the utility of the MT-REXEE method in increasing conformational sampling and predicting physically realistic complex conformations, but briefly we will review how restriction of loop motions provides a physical explanation for the observed differences in FE profile for the G107M FabB variant (Figure 16). Structural analyses of the trajectories suggest that the reason for this destabilization is that the G107M mutation restricts the accessible binding pocket for substrates longer than 8 carbons.
Figure 5: (A) The G107M mutation (red) blocks access to a region of pocket A which destabilizes chain binding to this pocket, but has little direct affect on pocket B. We show centroids for the C10 to C12 transformation for the chain in pocket A (brown) and pocket B (orange). To assess the effect of this mutation on the binding complex, we compare the binding free energy differences by chain length for the WT and G107M variant in pocket A (B) and pocket B (C). (D) To highlight the difference in stability between the two pockets we then compare the difference in individual relative binding free energies between pocket A and B for all chain lengths for the WT and G107M variant. There is a clear bias in the mutant FabB variant against the growth of the acyl chain from 6 to 8 and from 8 to 10 carbons.
There is a loop consisting of residues 105–110 in FabB which has two conformations. This loop reorientation occurs only for substrate C8 or longer in the WT FabB simulations. The G107M mutations prevent loop reorientation, which restricts the conformational space accessible by the acyl chain. Additionally, only in the reoriented conformation is this loop able to form stable non-bonded interactions with acyl chain carbons 7–9 which could be responsible for the destabilization of growing in C8 and C10 acyl substrates to FabB in the G107M variant. Once the chain has grown to 10 carbons we no longer see this additional cost for further extending the chain suggesting that the penalty is specifically for growing C8 and C10 chains which would not longer be included for growing in C12 or larger acyl chains. Further experimentation is necessary to validate this mechanism as being the primary contribution to this thermodynamic effect, but we have demonstrated the ability of the MT-REXEE to provide both thermodynamic quantification as well as a physical explanation for experimentally observed protein behavior in highly flexible systems.
5 Discussion and Conclusions
The MT-REXEE method increases conformational sampling for sets of alchemical transformations in which prohibitively high free energy barriers are present at only a subset of the states of interest. We use the FabB—acyl complex to demonstrate the potential of MT-REXEE to increase conformational sampling. In the FabB—acyl complex, two distinct pockets can be occupied by the acyl chain. Pocket B was inaccessible using standard MD simulations initiated in pocket A for any chain length longer than 4 carbons. MT-REXEE allowed for significantly enhanced sampling between the two pockets within a single trajectory for each transformation. This enhanced sampling has then been leveraged to quantify the difference in binding free energy surfaces between pocket A and pocket B for the G107M FabB variant, which has a known effect of preventing medium chain synthesis through pocket A.
This work also highlights the new exchange method, random range (RR), in which conformational exchanges are allowed at any frame within a given percentage of the trajectory rather than the last frame of an iteration. This improvement removes the necessity of including redundant or up-weighted end states which was previously necessary. We verified that this novel exchange method does not introduce additional error into the calculations through comparison of binding free energy estimates to MUP1 using both EE and MT-REXEE using the previous exhaustive method. All three methods produce statistically equivalent binding FE estimates, but we observe significantly faster convergence using MT-REXEE with the RR method. The RR procedure significantly reduces the round trip times in alchemical space while also increasing relative sampling of intermediate $\lambda$ states with the same trajectory length. Decreasing the round trip time increases the efficiency of the MT-REXEE method, which is of growing importance when simulating larger systems like that of the FabB–acyl complex. There is a trade-off when using the RR method in which an average of $\sim$25% of each iteration’s sampling (for the settings used in this study) is not included in the trajectories leading to conformational swaps. This sampling is still included in FE and trajectory analysis and is thus not lost, but will reduce the degree of decorrelation of the pre-swap configurations because of decreased total sampling time along continuous trajectories.
For the FabB–acyl complex, we simulate the covalent intermediate state. This state was selected to simplify the test system, compared to the FabB–acyl-ACP complex. Much like the formation of covalent bonds in intermediate states of enzymatic processes, covalent drugs initiate the formation of covalent bonds with their protein targets. Alchemical free energy estimates for reversible covalent drug compounds have been successfully computed, but often for relatively rigid ligands 4. The conformational sampling problem caused by increased conformational flexibility in the ligand is exacerbated by the anchor provided by the covalent bond which limits potential ligand and protein target rearrangements. The application of MT-REXEE to the FabB–acyl system demonstrates that converged estimates for the binding free energy for flexible covalent ligands can be achieved with reasonable computational resources without needing to identify or define collective variables.
Supplementary Information
Non-additive nature of $\Delta A_{ij,A}$ and $\Delta A_{ij,B}$
The Gibbs FE differences are not directly additive $\Delta G_{ij}\neq\Delta G_{ij,A}+\Delta G_{ij,B}$ as this would assume equal contributions from both configurational states. To get a direct expression for the free energy difference between states $i$ and $j$ regardless of configurational state ($\Delta G_{ij}$) in terms of $\Delta G_{ij,A}$ and $\Delta G_{ij,B}$ we start by rearranging equation 2 to give us equation 9 and 10. For simplicity’s sake, we utilize the variables $A_{i}$, $A_{j}$, $B_{i}$, and $B_{j}$ as defined in equations 9 and 10.
$\frac{A_{j}}{A_{i}}=\frac{\int_{\Gamma_{j,A}}e^{-\frac{U_{j}(\vec{q})}{k_{B}T}}d\vec{q}}{\int_{\Gamma_{i,A}}e^{-\frac{U_{i}(\vec{q})}{k_{B}T}}d\vec{q}}=e^{\frac{-\Delta A_{ij,A}}{k_{B}T}}$ (9)
$\frac{B_{j}}{B_{i}}=\frac{\int_{\Gamma_{j,B}}e^{-\frac{U_{j}(\vec{q})}{k_{B}T}}d\vec{q}}{\int_{\Gamma_{i,B}}e^{-\frac{U_{i}(\vec{q})}{k_{B}T}}d\vec{q}}=e^{\frac{-\Delta A_{ij,B}}{k_{B}T}}$ (10)
With this reduced representation we can rewrite equation 1 as $\Delta G_{ij}=-k_{B}T\ln{\frac{A_{i}+B_{i}}{A_{j}+B_{j}}}$. When we substitute in the values for Equations 9 and 10 we get
$\Delta G_{ij}=-k_{B}T\ln{\frac{e^{\frac{-\Delta A_{ij,A}}{k_{B}T}}+e^{\frac{-\Delta A_{ij,B}}{k_{B}T}}}{1+\frac{A_{i}}{B_{i}}}}$ (11)
which can finally be reduced to Equation 12 in which the final term represents the adjustment that must be made to account for uneven relative state weighting.
$\Delta G_{ij}=\Delta G_{ij,A}+\Delta G_{ij,B}+k_{B}T\ln{(1+\frac{\int_{\Gamma_{i,A}}e^{-\frac{U_{i}(\vec{q})}{k_{B}T}}d\vec{q}}{\int_{\Gamma_{i,B}}e^{-\frac{U_{i}(\vec{q})}{k_{B}T}}d\vec{q}})}$ (12)
We can directly compute $\Delta G_{ij}$ from Equation 1 with all trajectory frames rather than performing the separation in configurational space needed to calculate $\Delta G_{ij,A}$ and $\Delta G_{ij,B}$.
Figure 6: This diagram shows how we can discretize trajectories based on whether they are sampling from conformational space A(orange) or B(pink). Over the course of a single iteration the system will perform Monte Carlo transitions between $\lambda$ states which is represented by movement along the y-axis. The system will also freely move in configurational space between space A and B shown by changes in vertical line color. In analysis we can separate the trajectory into frames which sample in configurational space A and B (dashed red lines). We concatenate the energies for all $\lambda$ states sampled in configurational space A and B and then perform separate MBAR analysis to obtain the free energy estimates.
Figure 7: The separation of the free energy into conformational and alchemical free energy produces three distinct free energy cycles. One cycle follows the blue and pink arrows computing the relative free energy between Cn to Cn+2. The other two cycles follow the pink and either light or dark orange arrows computing the relative free energy for Cn between pockets A and B and separately Cn+2 and pockets A or B.
Figure 8: (A) We highlight the location of our reference residue THR300 (blue) position relative to the chain conformation in pocket A (orange) compared to pocket B (purple). (B) We plot the distance distributions in a set of trajectories with chain lengths ranging from 2 to 16 carbons and known to be in pocket A or B. The independent peaks demonstrate that this is an ideal metric for determining quantitatively defining the two pockets.
Figure 9: We compare the Wang-Landau weight equilibration times for all MUP1 small molecule ligand transformations performed both in solvent (A) and in the protein complex (B). The weight equilibration times for the MT-REXEE method with the Random Range equilibration method and no redundant end-states results in either statistically equivalent or lower equilibration times.
Figure 10: Here we use the absolute value of the difference in FE between using the first and last 50% of the trajectory following Wang-Landau weight equilibration. If the estimates are closer together at 50% we can conclude that the free energy estimates are converging more quickly which is likely a result of increased conformational sampling. (A) In the solvent simulations with the MUP1 ligands we see no real significant difference in the rate of convergence between methods. However in the protein complex simulations (B), we see a significant decrease in the difference between the forward and reverse estimates which indicates faster conformational sampling with MT-REXEE. (C) We also use the standard error between replicates as a measure of convergence for the free energy simulations which shows a significant decrease for the exhaustive method and an even more drastic decrease for MT-REXEE RR method.
Figure 11: (A) The round trip time is the time it takes for a conformation to swap through all parallel simulations and return to the original $\lambda$ state. The round trip times are divided into the first and last half of the trajectories. We see a significant increase in the round trip time in the latter half of the simulation. (B) We compare the $\lambda$ ratio for all $\lambda$ states in each transformation in the first and second half of the trajectory. This is computed as $R=\frac{N_{i}}{N_{avg}}$ where $N_{i}$ is the number of frames sampled at a given $\lambda$ states, and $N_{avg}$ is the average frames sampled for all $\lambda$ states. In an ideal situation R = 1 because every state is sampled evenly. A value of R¡1 means the state is under sampled and R¿1 means the state is over-sampled. We observe a corresponding shift in the observed $\lambda$ ratio in the latter half of the simulation. This is possibly because the sampled conformational space is deviating significantly from the space sampled as the weights were equilibrating.
Figure 12: As the solvent reference state here we utilize the ACP with a sequestered acyl chain. We show centroid structures from the MT-REXEE simulations of the solvent state for C2-4 (A) and the C14-16 (B) states.
Figure 13: We measure the configurational overlap between $\lambda$ states in the ACP solvent simulations for transformations (A) A-B, (B) B-C, (C) C-D, (D) D-E, (E) E-F, (F) F-G, (G) G-H. We optimized the $\lambda$ state spacing in order to ensure off-diagonal values of at least 0.05.
Figure 14: We measure the configurational overlap between $\lambda$ states in the FabB complex simulations for transformations (A) A-B, (B) B-C, (C) C-D, (D) D-E, (E) E-F, (F) F-G, (G) G-H. We optimized the $\lambda$ state spacing in order to ensure off-diagonal values of at least 0.05.
Figure 15: We present an example schematic for computing the cumulative FE difference to a single reference point as used in figure 4. In this example we compute the relative free energy difference between a C2 chain in pocket A and a C6 chain in pocket B. We add the configurational FE difference between pockets A and B for a C2 chain ($\Delta G_{conf,0}$) to the alchemical free energy difference of performing the C2 to C4, C4 to C6, and C6 to C8 transformations in pocket B ($\Delta G_{0,B}$, $\Delta G_{1,B}$, $\Delta G_{2,B}$ respectively).
Figure 16: The G107M mutations is observed to add a free energy cost for growing the acyl chain into pocket A for C6-8 and C8-10 transformations using MT-REXEE simulations. One physical explanation is that the restriction of conformational space means the blue C12 conformation is accessible only in the WT while the C12 conformation in pink is the primary accessible conformation in the G107M FabB variant. This chain rearrangement coincides with the movement of a loop consisting of residues 105-110. This loop rearrangement is only possible in the WT and not the G107M variant and these loop residues form non-bonded interactions with carbons 7–10 on the acyl chain only in the WT conformation.